In 10 minutes, a heart can beat 700 times. At this rate, in how many minutes will a heart beat 140 times? At what rate can a hear beat?


1
Expert's answer
2021-09-13T17:28:41-0400
"rate=\dfrac{N}{t}=\dfrac{700\ times}{10\ min}=70\ times/min"

"t=\dfrac{N}{rate}=\dfrac{140\ times}{70\ times/min}=2\ min"

2 minutes


A hear can beat at rate 70 times/min.